# On extension modules of finite homological dimension

### v1

@@ 5. On Finite Homological dimension of Ext modules over CM rings
We may assume that $R$ is complete and therefore $R$ admits a canonical module $\omega$. By [5, Corollary 2.1.4], $\operatorname{grade}(M)=t-s$, so $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $i<t-s$ and $\operatorname{Ext}_{R}^{t-s}(M,R)\not=0$. Then by Corollary 5.7, $\operatorname{H-dim}_{R}(M)<\infty$ and (1) it is proved.
On the other hand, since $\operatorname{Ext}_{R}^{i}(M,R)=0$ for $i>t-r$, one can see by assumption that $\operatorname{G-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $i\geq 0$. Thus, given $i\geq 0$, by Corollary 5.10,
$\operatorname{Ext}_{R}^{i}(M,R)\otimes_{R}\omega\cong\operatorname{Ext}_{R}^{i}(M,\omega).$
It follows that the modules $\operatorname{Ext}^{i}_{R}(M,R)$ and $\operatorname{Ext}^{i}_{R}(M,\omega)$ have the same support and so $\dim_{R}(\operatorname{Ext}_{R}^{i}(M,R))=\dim_{R}(\operatorname{Ext}_{R}^{i}(M,\omega))$. In this way, item (2) follows from [33, Lemma 1.9 a) and c)].
Item (3) follows from Corollary 5.10 and [13, Theorem 3.1], and item (4) follows from (3) and [5, Theorem 3.2.10(b)]. ∎
Corollary 5.12.
Assume $R$ is Cohen-Macaulay, and let $M$ be an $R$-module of grade $g$. If $M$ is Cohen-Macaulay and $\operatorname{H-dim}(\operatorname{Ext}_{R}^{g}(M,R))<\infty,$ then $M$ is $\operatorname{H}$-perfect.
Proof.
By [5, Corollary 2.1.4], $g=t-s$ with $s=\dim_{R}(M)$, so by assumption $\operatorname{G-dim}_{R}(\operatorname{Ext}_{R}^{t-s}(M,R))<\infty$. Since $M$ is Cohen-Macaulay, it follows from Corollary 5.11(1), that $\operatorname{H-dim}_{R}(M)<\infty$. It follows from the formula Proposition 2.3(2) that $\operatorname{H-dim}_{R}(M)=t-s=g$, i.e., $M$ is $\operatorname{H}$-perfect. ∎
@@
In [16, Theorem 2 and 2’], Foxby showed that if $R$ is Cohen-Macaulay, then an $R$-module $M$ of grade $g$ is $\operatorname{H}$-perfect if and only if $M$ is Cohen-Macaulay and $\operatorname{Ext}_{R}^{g}(M,R)$ is $\operatorname{H}$-perfect. Note that Corollary 5.12 generalizes the converse of that result.
In view of [18, Theorem 3.3], it should be noted that Corollary 5.11(3) generalizes the Ghosh-Puthenpurakal type formula [18, Theorem 3.5] for Cohen-Macaulay rings and answers negatively [18, Question 3.9] when we are over Gorenstein rings.
Corollary 5.14.
Assume $R$ is Gorenstein. For any $R$-module $M$ one has
$\operatorname{r}(M)=\mu(\operatorname{Ext}_{R}^{\operatorname{G-dim}_{R}(M)}(M,R)).$
Although most of the next result is known, our results (including Theorem 5.6) show that such different characterizations of the Gorenstein property of a Cohen-Macaulay ring come from the same place.
Corollary 5.15.
Assume $R$ is Cohen-Macaulay and let $\omega$ be its canonical module. We let $\_^{\dagger}$ stands for $\operatorname{Hom}_{R}(\_,\omega)$. The following are equivalent:
  1. (1)
    $R$ is Gorenstein.
  2. (2)
    $\operatorname{Gid}_{R}(k)<\infty$.
  3. (3)
    $\operatorname{Gid}(R)<\infty$.
  4. (4)
    $\operatorname{G-dim}_{R}(\omega)<\infty$.
  5. (5)
    $\operatorname{G-dim}_{R}(\omega^{*})<\infty$.
    Now consider the following statements provided that $\operatorname{Ext}^{i}_{R}(\omega,R)=0$ for all $1\leq i\leq t$.
  6. (6)
    $\operatorname{Gid}_{R}((\omega^{*})^{\dagger})<\infty$.
  7. (7)
    $\operatorname{Gid}_{R}(\omega^{*})<\infty$.
  8. (8)
    $\operatorname{G-dim}_{R}((\omega^{*})^{\dagger})<\infty$.
Then $\rm(1)-(5)$ are mutually equivalent, and each of them is equivalent to $\rm(6)$. In addition, $(6)\Rightarrow(7)\Leftrightarrow(8)$.
Proof.
The equivalences $(1)\Leftrightarrow(2)$ and $(3)\Leftrightarrow(4)$ follow directly from Corollary 4.2. Also, Theorem 5.6 asserts immediately that $(3)\Leftrightarrow(5)$. The equivalence $(1)\Leftrightarrow(3)$ is proved in [22, Theorem 2.1].
Now, note that from the vanishing hypothesis and [21, Theorem 2.1(i)] we see that $\omega^{*}$ is maximal Cohen-Macaulay. Therefore $(5)\Leftrightarrow(6)$ and $(7)\Leftrightarrow(8)$ follow from Corollary 4.2. Since $(1)\Leftrightarrow(6)$, it is clear that $(6)\Rightarrow(7)$. ∎
@@ 6. Some questions
In this section, we discuss some natural questions motivated by the results established earlier. Our primary question stems from the challenge of extending Corollaries 5.7 and 5.8 to the non-Cohen-Macaulay setting, removing the requirement that $R$ be a Cohen-Macaulay ring.
Question 6.1.
@@
    If $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$ and $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $1\leq i\leq t$, then $\operatorname{H-dim}_{R}(M)=0$.
The following proposition partially answers Question 6.1 under the assumption that $\operatorname{H\text{-}dim}_{R}=\operatorname{pd}_{R}$.
@@
    $M$ has finite Gorenstein dimension and $s=t-0pt(M)$
@@
In both cases, the implication ()($\Rightarrow$)( ⇒ ) follows from Proposition 3.4. Thus, it remains to prove the implication ()($\Leftarrow$)( ⇐ ).
(1) Suppose that $\operatorname{pd}_{R}(N)<\infty$. By Theorem 2.1 and Lemma 5.4, $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(\Omega^{t}(M),N))<\infty$. Moreover, the vanishing condition shows that $\operatorname{Ext}_{R}^{i}(\Omega^{t}(M),R)=0$ for all $1\leq i\leq t$. Therefore, by [12, Proposition 19], $\operatorname{pd}_{R}(\Omega^{t}(M)^{\ast})<\infty$. Then by [12, Proposition 13], $\Omega^{t}(M)$ is free. Thus, $\operatorname{pd}_{R}(M)<\infty$.
@@
Remark 6.3.
It should be noticed that in Proposition 6.2 the implication $\Rightarrow$ only requires that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for $0\leq i\leq t-\operatorname{depth}(M)$.
For $N=R$, an answer to Question 6.1(2) is given by [12, Proposition 13]. Next, we present an answer to Question 6.1(1) in the following proposition, which follows from and strengthens [12, Proposition 13] for projective dimension and Gorenstein dimension. For that, we say that an $R$-module $M$ satisfies $(\widetilde{S}_{n})$ if $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\min\{n,0pt_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\}$ for all prime ideals $\mathfrak{p}$ of $R$.
@@
Let $s=t-n$. Then by Theorem 2.1 and Lemma 5.4, $\operatorname{H-dim}_{R}(\Omega^{s}(M)^{\ast})<\infty$. Since $\operatorname{H-dim}_{R}(M^{\ast})<\infty$ and $M$ satisfies $(\widetilde{S}_{n})$, an application of the depth lemma shows that $\Omega^{s}(M)$ satisfies $(\widetilde{S}_{t})$. It follows from [12, Proposition 13] that $\operatorname{H-dim}_{R}(\Omega^{s}(M))=0$. Hence, by Proposition 2.3(3), $\operatorname{H-dim}_{R}(M)\leq s=t-n$. ∎
Remark 6.5.
It is important to mention Proposition 6.4 is significant only when $\operatorname{grade}(M)=0$ or $n=0$. The reason of this is that there are no non-zero $R$-modules $N$ with positive grade satisfying $(\widetilde{S}_{1})$. The proof of this fact is as follows. Suppose that $N$ satisfies $(\widetilde{S}_{1})$ and that $N^{\ast}=0$. Then, $\operatorname{pd}_{R}(\operatorname{Tr}_{R}(N))\leq 1$. As $N$ satisfies $(\widetilde{S}_{1})$, by [36, Theorem 5.8], we have that $\operatorname{Ext}_{R}^{1}(\operatorname{Tr}_{R}(N),R)=0$. Thus, $\operatorname{Tr}_{R}(N)$ is free and hence $N$ as well. But since $N^{\ast}=0,$ then $N=0$.
@@
Note that Proposition 6.4 and Corollary 6.6 give some answers to [28, Question 4.11]. As an application of Corollary 6.6, we can improve [19, Proposition 3.3] as follows:
@@
From the assumption we see that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $0\leq i\leq t$. Then by Corollary 6.6, $\operatorname{H-dim}_{R}(M)<\infty$. Since $\operatorname{H-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,R)\not=0\}$ (see, for instance [26, Lemma 2.3(c)] together with [2, Theorem 1.4]) and $\operatorname{H-dim}_{R}(M)\leq t$, we see from the vanishing condition that $\operatorname{H-dim}_{R}(M)=h$. ∎
The reader can note that Corollaries 5.7 and 5.8, which motivated Question 6.1, were obtained using Theorem 5.6, which was derived from Theorem 5.5. Thus, in view of the Foxby equivalence ([6, Theorem 3.3.2]), trying to answer these questions, we pose the following question:
Question 6.8.
@@
Question 6.9.
Let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t$. Then does the equivalence
@@
One can note that the implication ()($\Leftarrow$)( ⇐ ) holds by Theorem 3.6. On the other hand, the other implication has a positive answer if $M$ is maximal Cohen-Macaulay. Indeed, by the Bass conjecture, $R$ is Cohen-Macaulay, and we may assume (passing to completion if necessary) that $R$ admits a canonical module. Then $M\cong\omega^{n}$ for some $n>0$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(\omega,N))<\infty$, obtaining thus the desired conclusion from Proposition 5.1(3).
Based on the blend between the Gorenstein dimensions, the canonical module, and its duals in Corollary 5.15, we close the paper by asking the following.
Question 6.10.
@@
Assume $R$ is Cohen-Macaulay with canonical module $\omega$, and let $\_^{\dagger}$ stands for $\operatorname{Hom}_{R}(\_,\omega)$. Does any of the conditions
  1. (1)
    $\operatorname{Gid}_{R}(\omega^{*})<\infty$;
  2. (2)
    $\operatorname{G-dim}_{R}((\omega^{*})^{\dagger})<\infty$.
imply that $R$ must be Gorenstein?
Acknowledgment. This work was completed while the third author was visiting the University of Nebraska-Lincoln. He is grateful for the hospitality. He also thanks Roger and Sylvia Wiegand, Ryo Takahashi, Hailong Dao, Paulo Martins, Nawaj KC, Zachary Nason, Anna Brosowsky, Andrew Soto Levins and Kesavan Mohana Sundaram for some discussions about this work. In addition, he also thanks to Lars W. Christensen and Henrik Holm for their response to a question about completion and Gorenstein injective dimension. Additionally, the authors are grateful to Ryo Takahashi, Hailong Dao and Nawaj KC for suggesting some examples, and to Dipankar Ghosh for correcting some typos.

### v2

@@ 4. A duality for Ext modules with finite homological dimension
We may assume that $R$ is complete and therefore $R$ admits a canonical module $\omega$. By [5, Corollary 2.1.4], $\operatorname{grade}(M)=t-s$, so $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all i<tsi<t-s and $\operatorname{Ext}_{R}^{t-s}(M,R)\not=0$. Then by Corollary 4.7, $\operatorname{H-dim}_{R}(M)<\infty$.
On the other hand, since $\operatorname{Ext}_{R}^{i}(M,R)=0$ for i>tri>t-r, one can see by assumption that $\operatorname{G-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $i\geq 0$. Thus, given $i\geq 0$, by Corollary 4.9, $\operatorname{Ext}_{R}^{i}(M,R)\otimes_{R}\omega\cong\operatorname{Ext}_{R}^{i}(M,\omega).$ It follows that the modules $\operatorname{Ext}^{i}_{R}(M,R)$ and $\operatorname{Ext}^{i}_{R}(M,\omega)$ have the same support and so $\dim_{R}(\operatorname{Ext}_{R}^{i}(M,R))=\dim_{R}(\operatorname{Ext}_{R}^{i}(M,\omega))$. In this way, the last assertion follows from [29, Lemma 1.9 a) and c)]. ∎
@@
Ghosh and Puthenpurakal proved in [15, Theorem 3.5] that if $M$ is a finitely generated $R$-module that is $\operatorname{G}$-perfect (i.e., $\operatorname{grade}(M)=\operatorname{G\text{-}dim}_{R}(M)$), then $\operatorname{r}(M)=\mu(\operatorname{Ext}^{t-r}_{R}(M,R))\operatorname{r}(R)$. A more general formula was later obtained in [32, Theorem (A)]. As an application of Corollary 4.9, one can prove the Ghosh–Puthenpurakal formula to finitely generated $R$-modules $M$ over a Cohen–Macaulay ring RRsatisfying $\operatorname{G\text{-}dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $t-s\leq i\leq t-r$. Indeed, as in the proof of Corollary 4.10, this finiteness condition holds for all $i\geq 0$. Hence Corollary 4.9 yields $\mu(\operatorname{Ext}_{R}^{i}(M,\omega))=\mu(\operatorname{Ext}_{R}^{i}(M,R))\operatorname{r}(R)$, and the formula follows from [12, Theorem 3.1].
@@ 5. Some questions
In this section, we discuss some natural questions motivated by the results established earlier. Our primary question stems from the challenge of extending Corollary 4.7 to the non-Cohen-Macaulay setting.
Question 5.1.
@@
    If $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$ and $\operatorname{Ext}_{R}^{1\leq i\leq t}(M,N)=0$, then $\operatorname{H-dim}_{R}(M)=~0$.
The following proposition partially answers Question 5.1 when $\operatorname{H\text{-}dim}_{R}=\operatorname{pd}_{R}$.
@@
    $M$ has finite Gorenstein dimension equal to $s$.
@@
In both cases, the implication $(\Rightarrow)$ follows from Corollary 2.3. Thus, it remains to prove the implication $(\Leftarrow)$.
(1) Suppose that $\operatorname{pd}_{R}(N)<\infty$. By Lemma 4.4, we have $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(\Omega^{t}(M),N))<\infty$. Moreover, the vanishing condition shows that $\operatorname{Ext}_{R}^{i}(\Omega^{t}(M),R)=0$ for all $1\leq i\leq t$. Therefore, by [11, Proposition 19], $\operatorname{pd}_{R}(\Omega^{t}(M)^{\ast})<\infty$. Then by [11, Proposition 13], $\Omega^{t}(M)$ is free. Thus, $\operatorname{pd}_{R}(M)<\infty$.
@@
For $N=R$, an answer to Question 5.1(2) is given by [11, Proposition 13]. Next, we present an answer to Question 5.1(1) that strengthens this proposition. For that, we say that an $R$-module $M$ satisfies $(\widetilde{S}_{n})$ if $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\min\{n,0pt_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\}$ for all prime ideals $\mathfrak{p}$ of $R$.
@@
Let s=tns=t-n. Then by Lemma 4.4, $\operatorname{H-dim}_{R}(\Omega^{s}(M)^{\ast})<\infty$. Since $\operatorname{H-dim}_{R}(M^{\ast})<\infty$ and $M$ satisfies $(\widetilde{S}_{n})$, an application of the depth lemma shows that $\Omega^{s}(M)$ satisfies $(\widetilde{S}_{t})$. It follows from [11, Proposition 13] that $\operatorname{H-dim}_{R}(\Omega^{s}(M))=0$. Hence, by [4, Theorem 8.7(3)], $\operatorname{H-dim}_{R}(M)\leq s=t-n$. ∎
@@
Note that Proposition 5.3 and Corollary 5.4 give some answers to [24, Question 4.11]. As an application of Corollary 5.4, we can improve [16, Proposition 3.3].
@@
From the assumption we see that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $0\leq i\leq t$. Then by Corollary 5.4, $\operatorname{H-dim}_{R}(M)<\infty$. Since $\operatorname{H-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,R)\not=0\}$ (see, for instance [22, Lemma 2.3(c)] together with [2, Theorem 1.4]) and $\operatorname{H-dim}_{R}(M)\leq t$, we see from the vanishing condition that $\operatorname{H-dim}_{R}(M)=~h$. ∎
The reader can note that Corollary 4.7, which motivated Question 5.1, was obtained using Theorem 4.6, which was derived from Theorem 4.5. Thus, in view of the Foxby equivalence ([6, Theorem 3.3.2]), we pose the following question:
Question 5.6.
@@
Question 5.7.
Let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t$ (or all $i\geq 0$). Then does the equivalence
@@
Note that the implication $(\Leftarrow)$ holds by Theorem 2.5. On the other hand, the other implication has a positive answer if $M$ is maximal Cohen-Macaulay. Indeed, by the Bass conjecture, $R$ is Cohen-Macaulay, and we may assume (passing to completion if necessary) that $R$ admits a canonical module. Then $M\cong\omega^{n}$ for some $n>0$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(\omega,N))<\infty$, obtaining thus the desired conclusion from Proposition 4.1(3).
Motivated by Corollary 3.3 and [24, Corollary 5.13], we close this section by asking the following.
Question 5.8.
@@
Assume $R$ is Cohen-Macaulay with canonical module $\omega$, and let $^{\dagger}$ stand for $\operatorname{Hom}_{R}(\_,\omega)$. Does any of the conditions $\operatorname{Gid}_{R}(\omega^{*})<\infty$ or $\operatorname{G-dim}_{R}((\omega^{*})^{\dagger})<\infty$ imply that $R$ must be Gorenstein?
Addendum
After the first version of this paper, recently Questions 5.1 and 5.7 were affirmatively answered by Kaito Kimura in [20]. The technique in his paper differs from the one we used in Theorem 4.7.
6. Examples
We present some examples concerning Ext modules with finite homological dimension.
Example 6.1.
Assume $R$ is a local ring and let $\underline{x}=x_{1},\ldots,x_{n}$ be a regular sequence. Let $N$ be an $R$-module such that $\underline{x}N=0$. Set $M=R/(\underline{x})$. Since the minimal free resolution of $M$ is the Koszul complex of the sequence $\underline{x}$, we have
$\operatorname{Ext}^{i}_{R}(L,X)\cong\operatorname{Ext}^{i}_{R}(R/(x_{1},\ldots,x_{i}),X)\cong X/(x_{1},\ldots,x_{i})X.$
Thus $\operatorname{H-dim}_{R}(X)<\infty$ if and only if $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(L,X))<\infty$ for all $0\leq i\leq n$. In particular, $\operatorname{H-dim}_{R}(\operatorname{Ext}^{i}_{R}(L,R))=i$, by the Auslander-Buchsbaum and Auslander-Bridger formulas.
An $R$-module $M$ is said to be $\operatorname{H-}$perfect if $\operatorname{H-dim}_{R}(M)<\infty$ and $\operatorname{grade}(M)=\operatorname{H-dim}_{R}(M)$.
Example 6.2.
Let $M_{1},\ldots,M_{n}$ be H-perfect $R$-modules of grade $g_{1},\ldots,g_{n}$, respectively. Then each $\operatorname{Ext}_{R}^{g_{i}}(M_{i},R)$ have $\operatorname{H-dim}$ equal $g_{i}$ by [3, Theorem 6.3]. Set $M=\bigoplus_{i=1}^{n}M_{i}$. Thus, for $j\not=g_{1},\ldots,g_{n}$, the $R$-module $\operatorname{Ext}_{R}^{j}(M,R)$ is zero, while that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{g_{i}}(M,R))=g_{i}$ for each $1\leq i\leq n$.
An $R$-module $M$ is $p$-spherical if $\operatorname{pd}_{R}(M)\leq p$ and $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq p-1$.
Example 6.3.
Based on the proof of [21, Proposition 6] we can construct spherical non-perfect modules as follows. We repeat the proof for convenience. Let $M$ be an $R$-module with positive grade $n$, and let $F_{\bullet}$ be a free resolution of $M$. Consider the truncation
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${F_{n-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{0}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{1}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{2}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle$$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${F_{n-1}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}}$${F_{n}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
is a free resolution of $N$ and
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${F_{n-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
is exact. This means that $N$ is $n$-spherical as well as $N^{*}$ is a $(n+1)$th syzygy of $M$ (in the sense of [22]), and also $\operatorname{Ext}^{n}_{R}(N,R)\cong M$. It follows that whenever $M$ is such that $\operatorname{pd}_{R}(M)>\operatorname{grade}(M)=n>0$, there exists a $n$-spherical non-perfect $R$-module NN such that $\operatorname{Ext}^{n}_{R}(N,R)\cong M$. In particular, if $\operatorname{H-dim}_{R}(M)<\infty$ (resp. $\operatorname{Hid}_{R}(M)<\infty)$ then $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{n}(N,R))<\infty$ (resp. $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{n}(N,R))<\infty$).
Example 6.4.
$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${R\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x}$${R\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{y}$${R\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x}$${R\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
so that $\operatorname{Ext}^{2i+1}_{R}(M,M)\cong M$ for all $i$, which means that $\operatorname{G-dim}_{R}(\operatorname{Ext}^{2i+1}_{R}(M,M))=0$ and $\operatorname{pd}_{R}(\operatorname{Ext}^{2i+1}_{R}(M,M))=\infty$.